### v2

# What Do the Circuits Mean? A Knowledge Edit View

Abstract
In the field of language model interpretability, circuit discovery is gaining popularity. Despite this, the true meaning of these circuits remains largely unanswered. We introduce a novel method to learn their meanings as a holistic object through the lens of knowledge editing. We extract circuits in the GPT-2 base model (Radford et al., 2019) for classification tasks related to syntax and model safety, and study their knowledge property via a model edit dataset containing hierarchical entities (Ge et al., 2024). We find that these circuits contain entity knowledge but resist new knowledge, demonstrating a "confirmation bias" behavior. Additionally, we examine the impact of circuit size, discovering that an ideal "theoretical circuit" where essential knowledge is concentrated likely incorporates more than 5% but less than 50% of the model’s parameters. We also assess the overlap between circuits from different datasets, finding moderate similarities. We proceed with analyzing the modular components of the circuits, finding that up to 60% of the circuits consist of layer normalization modules rather than attention or MLP modules, adding evidence to the ongoing debates regarding knowledge localization. In summary, our findings offer novel insights into the meanings of the circuits, and introduce directions for further interpretability and safety research of language models.
What Do the Circuits Mean? A Knowledge Edit View
Huaizhi Ge Columbia University hg2590@columbia.edu                        Frank Rudzicz Dalhousie University frank@dal.ca                        Zining Zhu Stevens Institute of Technology zzhu41@stevens.edu
1 Introduction
The interpretability of large language models (LLMs) has advanced significantly with the exploration of "circuits"—specific subsets of model parameters tailored for distinct tasks. Derived from specialized training on text datasets, these circuits offer profound insights into how LLMs encode and process information. What do the circuits mean? To address this question, an avenue of existing works applied ablation studies on designated components of the circuits (Conmy et al., 2023), which require laborious expert scrutiny. Another promising avenue involves verbalizing some parts of the networks, either using the language model head as a "lens" (nostalgebraist, 2020; Belrose et al., 2023) or an external model (Hernandez et al., 2022; Singh et al., 2023), but their approaches have limited applicability to automatically-extracted circuits.
We propose an approach to study the meanings of the automatically-extracted circuits as a holistic object (rather than its components), through the lens of knowledge editing.
We first investigate the storage locations of knowledge within the GPT-2 base model (Radford et al., 2019) using circuits derived from diverse text classification datasets. These circuits, taking up 50% of the model’s parameters, were adapted to generation tasks to facilitate knowledge editing tasks. Using fine-tuning as the edit method, we assess the knowledge editing capabilities of these circuits. The primary 50% circuit is presumed to contain pertinent knowledge, while the complementary 50% circuit is considered less integral to knowledge storage. We also find a "confirmation bias" behavior on the circuit level: The knowledge-intense circuits show stronger resistance to knowledge editing than the less knowledgeable complementary parts of the network.
Then, we explore the impact of reducing the circuit size. Initially working with circuits comprising 50% of the model’s parameters, we investigate circuits accounting for 5% of the parameters. This allow us to examine how circuit size influences their ability to retain and modify embedded knowledge. Our results suggest that an ideal "theoretical circuit" where essential knowledge is densely concentrated is likely larger than the 5% circuit and smaller than the 50% circuit.
Additionally, we investigate whether circuits from various datasets significantly overlap, despite their diverse origins. Using a heatmap to illustrate mask similarity across tasks, we observe that the extent of overlap between circuits varies moderately, suggesting nuanced similarities rather than straightforward congruence or divergence. Notably, circuits from datasets representing linguistic tasks and knowledge tasks show specific degrees of overlap, highlighting the complex ways these circuits might resemble each other.
Lastly, our study examines the roles of different types of layers within circuits, revealing that LayerNorm constitutes a large share of the circuits. Conversely, attention and multi-layer perceptron (MLP) layers hold smaller shares. This distribution challenges traditional views on knowledge storage in neural networks (which suggest that MLP layers may be vital knowledge repositories). The prominence of layer normalization layers underscores their potential impact on network stability and performance, highlighting the need for deeper investigation into how these layers influence the overall efficacy of LLMs.
In summary, we propose a perspective that can provide rich insights into the operational dynamics of circuits within LLMs. By examining the effects of circuit size, the overlap between circuits from different datasets, and the distinct roles of various layers, our study enhances the understanding of how neural networks manage and modify information. These findings enable future explorations into the precise manipulation of circuits, aiming to optimize model performance and interpretability in increasingly complex applications.
2 Related work
Circuit analysis
Circuit extraction tries to find a minimal subnetwork that represents the behavior of the full network (computation graph), where the components can be explainable (Olah et al., 2020; Elhage et al., 2021). This extracted circuit should be faithful Hanna et al. (2024): if the non-circuit graph edges outside of the circuit are removed, the model’s performance should stay the same. Each component of the circuit (the vertex of the graph) is usually a module Conmy et al. (2023); Wang et al. (2022) or a causal variable that supports interventions Vig et al. (2020); Geiger et al. (2021). Here, we focus on the parameter level, but our approach can be extended to other granularity levels without loss of generality.
Model editing
Various methods modify knowledge in large language models. Zhu et al. (2020) introduced a task for selective updates of factual knowledge in Transformer models, using fine-tuning as a benchmark. Rank-One Model Editing (ROME) by Meng et al. (2022a) adjusted feed-forward weights to change specific factual associations. MEMIT (Meng et al., 2022b) enabled integrating multiple memories into a language model. Model Editor Networks with Gradient Decomposition (MEND) by Mitchell et al. (2021) utilized a single targeted input-output pair for rapid, localized changes in a pre-trained model’s behavior. Other influential approaches include targeting specific knowledge neurons (Dai et al., 2021) and using hyper-networks (De Cao et al., 2021). These techniques have proven to be highly effective in editing knowledge within LLMs.
3 Methods
Circuit extraction
We adopt a differentiable masking technique implemented in Yu et al. (2024). Similar to the differentiable masking literature (De Cao et al., 2022; Csordás et al., 2020), we set up learning objectives to train a mask parameter $m_{p}\in[0,1]$ for each parameter $p\in\mathcal{P}$, where $\mathcal{P}$ describes the collection of all trainable parameters in the full model. Unlike De Cao et al. (2022), we use a combination of the following objectives:
  1. 1.
    Faithfulness loss $L_{F}=\frac{1}{N}\sum_{j}\hat{y}_{j}\textrm{log}\frac{1}{\hat{l}_{j}}$, which is a cross-entropy loss. This objective encourages the logits of the circuit $\hat{l}_{j}$ to predict the results predicted by the original model $\hat{y}_{j}$. This loss is averaged across the $N$ training data samples.
  2. 2.
    Sparseness loss $L_{S}=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}\sigma(m_{p})$, where $m_{p}$ is the mask for the model parameter $p$. This approximately counts the density of the nonzero masks (with a sigmoid function $\sigma(\cdot)$ that smooths the distribution). This encourages the circuit to be sparse.
Following De Cao et al. (2022), we use the hard-concrete distribution of Louizos et al. (2018) to sparsify the masks. Following Yu et al. (2024), we use a straight-through estimator (Bengio et al., 2013) to make the forward pass (apply the mask) trainable. We train until the sparseness reaches a pre-defined threshold — we experimented with 5% and 50% respectively, and their effects will be presented in subsequent sections.
Circuit-aware model edit
In the circuit extraction step, we identified a circuit $\mathcal{C}\subseteq\mathcal{P}$. To study how circuits are affected by the model’s knowledge, we propose a novel analysis technique, circuit-aware model editing. We employ fine-tuning, adapted from the EasyEdit repository (Wang et al., 2023), to modify the knowledge embedded in the model. The training objective is defined as:
$L_{C}=-\frac{1}{N}\sum_{i}\left[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right],$ (1)
where $N$ is the number of samples, $y_{i}$ is the true label for the $i$-th sample (0 or 1), and $p_{i}$ is the predicted probability of the $i$-th sample being in class 1. This objective represents the traditional Cross Entropy Loss. We use adjusted Stochastic Gradient Descent (SGD) for optimization, updating the parameters as follows:
$\theta_{t+1}=\theta_{t}-\eta\xi\nabla_{\theta}L_{C}(\theta_{t};x_{i},y_{i}),$ (2)
where $\theta_{t}$ are the parameters at iteration $t$, $\eta$ is the learning rate, and $\xi$ are the parameter masks, which are either 0 or 1. For parameters involved in the circuit $\mathcal{C}\subseteq\mathcal{P}$, the mask is set to 1; otherwise, it is set to 0. The term $\nabla_{\theta}L_{C}(\theta_{t};x_{i},y_{i})$ represents the gradient of the loss function $L_{C}$ with respect to the parameters $\theta_{t}$, computed using a single training example $(x_{i},y_{i})$ or a mini-batch of examples.
In this study, we specifically focus on editing the circuits within the model. Parameters not involved in these circuits are initially set to zero and are maintained at zero throughout the training process by setting their gradients to zero in each step, ensuring they remain unchanged.
4 Data
Model edit data
We use Hierarchy Data (Ge et al., 2024) to act as the model edit data. This dataset encompasses a series of both challenging incorrect facts, represented as $(s,r,o^{*})$, and their corresponding accurate facts, denoted as $(s,r,o)$. Here, $s$ signifies the subject and $o$ the object, both selected from the hierarchy chains. The relation $r$ consistently adopts the hypernym schema, emphasizing hierarchical connections. An example of this is "A British Shorthair is a kind of cat". We modify the objects of these facts, generating altered facts "A British Shorthair is a kind of dog" to test the efficacy of circuit edits. The Hierarchy Data dataset includes a corpus of 198 facts targeted for editing analysis. This structured approach facilitates explorations into the role of hierarchical relations in the adaptability and accuracy of language model editing processes.
Text classification data
To extract the circuit, we employ various datasets for text classification tasks. Specifically, we use eight syntax knowledge datasets from BLiMP (Warstadt et al., 2020), including ‘determiner noun agreement 1‘(dna1), ‘determiner noun agreement 2‘(dna2), ‘determiner noun agreement irregular 1‘(dnai1), ‘determiner noun agreement irregular 2‘(dnai2), ‘determiner noun agreement with adjective 1‘(dnawa1), ‘determiner noun agreement with adj 2‘(dnawa2), ‘determiner noun agreement with adj irregular 1‘(dnawai1), and ‘determiner noun agreement with adj irregular 2‘(dnawai2). Also we use five datasets from Model-Written Evaluations (Perez et al., 2022), including ‘risk averse‘(ra), ‘extraversion‘, ‘desire for being rated helpful, harmless & honest over actually being helpful, harmless & honest‘(hhh), ‘okay with using many resources‘(umr), and ‘desire for advancing technology to achieve goals‘(tech). Additionally, we established a text classification dataset using Hierarchy Data(h). For instance, we categorized the statement "A cat is a kind of animal" with a label of 1, indicating a true assertion, whereas the statement "A cat is a kind of plant" was labeled as 0, denoting a false assertion.
5 Evaluation of circuit extraction fidelity
To validate the accuracy and correctness of the extracted circuits, we evaluate their performance across specific classification tasks. In particular, we focus on two types of circuits: those containing 5% parameters and those containing 50% parameters.
Table 1 presents the task accuracy achieved by these identified circuits. The circuits were evaluated based on their ability to reproduce task outcomes. High accuracy scores indicate that the circuits not only preserve the task-relevant information but also reflect the underlying computational mechanisms of the original model.
The results show that both the 5% and 50% circuit subsets achieve high levels of accuracy, suggesting that the circuits have been correctly identified.
h circuit dna1 circuit dna2 circuit ra circuit e circuit
5% circuit 0.84 0.91 0.87 0.94 0.97
50% circuit 0.88 0.89 0.87 0.94 0.87
Table 1: Task accuracy of identified circuits
6 How knowledgeable are the LLM circuits?
6.1 Experiments
To investigate the amount of knowledge in the LLM circuits, we edit the knowledge within a specific circuit and compare its performance with that of its complementary circuit. Typically, individual circuits represent only a small percentage of the model’s total parameters. However, for a clearer demonstration of the amount of knowledge in the circuits, we analyze 50% circuits, ensuring that these had an equivalent number of parameters as their complementary circuits.
Initially, we use diverse text classification datasets to identify circuits within GPT-2 base model. These circuits constitute 50% of the entire model. Upon locating these circuits, we remove the classification head of the GPT-2 base model, transforming it into a generation model. Thus, we repurpose the identified circuits from text classification tasks for generation tasks. In these generation tasks, our aim is to perform knowledge editing. This involves altering the true target to a new target. For instance, given the prompt "A cat is a kind of," we would edit the true target "animal" to the new target "plant." Throughout this study, we employ fine-tuning techniques to modify the knowledge stored within these circuits.
In addition to the "50% circuit" derived from the text classification task, which we presume to be closely associated with the knowledge we aim to edit, we also refer to the remainder part of the model’s parameters "the complementary circuit".
6.2 The circuit vs complementary circuit
Figure 1: Pre-edit negative log probability of true target distribution of 50% circuits.
Even before applying any edits, we want to understand how the circuit and the complementary circuit encode entity-related knowledge, so we query the pre-edit negative log probabilities of the true targets. A lower values indicate a stronger grasp of the knowledge.
Figure 1 displays the results. The extracted circuits demonstrate significant lower (p-value < 0.05) negative log probabilities of the HierarchyData entities compared to their complementary circuits, indicating a superior understanding of the knowledge. Note that none of the circuits outperforms the full model, showing that the complementary circuits still capture some factual knowledge.
A noteworthy observation is that this effect generalizes across both the h circuit (which is derived from the same data as the HierarchyData entities) and the dna1 circuit (which is tasked with a distinct, syntax-focused dataset). This finding suggests that the model’s circuits for solving the syntax tasks may be relevant to the model’s mechanism for storing knowledge, thus providing a potential explanation for the "spill-over" effects (Sahak et al., 2023) — i.e., the effects of some datasets on the models may "spill over" to other apparently irrelevant datasets. We will study the overlap of circuits between different tasks in Section 8.
As a baseline, the probability of a random circuit is not significantly different from that of its complementary circuit, confirming that the aforementioned "circuit vs complementary" differences do not follow from randomness.
6.3 The complementary circuit is more susceptible to the model editing
After applying the circuit-aware knowledge editing, we query the negative log probabilities of the HierarchyData entities again, and the outcomes are presented in Figure 2.
Figure 2: Post-edit negative log probability of new target distribution of 50% circuits.
Our metric for assessing edit performance is the negative log probability of the new target, where lower values indicate better editing. The figure displays results for circuits extracted using the ‘determiner noun agreement 1’ (dna1) dataset of BLiMP, circuits extracted using hierarchy data (h) along with their complementary circuits, a randomly extracted circuit alongside its complementary circuit, and the full model circuit. Further details about other datasets are provided in Appendix A. Notably, the complementary 50% circuit consistently outperforms the 50% circuit in terms of editing for both the dna1 and h circuits. Furthermore, the performance of the random circuit is similar to that of its complementary counterpart. These indicate that the complementary circuits, which are less task-relevant, are more susceptible to knowledge edits. Yet, the full model demonstrates the best editing performance, indicating that the circuit might still synergize with the complementary networks upon the injection of new knowledge.
Similar to the pre-edit observations, the post-edit discrepancies between the circuits and their complements can be generalized to a linguistic task (dna1). This suggests that task-specific circuits might indicate the knowledge storage locations. In other words, altering the knowledge from these knowledge-rich circuits is more difficult than altering the knowledge from their complements. Interestingly, humans have similar "confirmation bias" behavior as well: it’s more difficult to change someone’s opinion once it’s formed.
Our findings here can also provide a novel perspective about neural network plasticity (Lyle et al., 2023), but a further in-depth inquiry into the nature of circuit-specific plasticity is left to future works.
7 How large are the ideal "theoretical circuits"?
In our previous experiments, we conducted experiments on circuits that comprised 50% of the total model parameters. To investigate potential disparities, we reduced our analysis to include circuits that accounted for 5% of the parameters. Figure 3 (a) illustrates the outcomes of editing the 5% vs. the 50% circuits, along with their corresponding complementary circuits, using HierarchyData. We continue to employ the negative log probability of the new targets to quantify the edit performance. As comparisons, Figure 3 (b) presents the negative log probability scores of the true targets before model edits. Detailed figures of 5% circuits derived from other datasets are presented in Appendix B.
Figure 3 (a) shows that both the 5% and the 50% circuits are more hard to edit than their corresponding complementary circuits. Notably, the 5% circuit displays a higher level of such "confirmation bias" than the 50% circuit. This suggests that the 5% circuit, rather than the 50% circuit, may be closer to an idealized "theoretical circuit" where essential knowledge is concentrated.
Now, if a theoretical circuit exists, would its size be smaller than 5% or between 5% and 50%?
As Figure 3 (b) shows, the 5% circuit shows the best capturing of knowledge, followed by the 50% circuit, because the 50% circuit may contain non-knowledgeable "noise" parameters that mask out the circuit’s ability to demonstrate knowledge. Also shown in this figure, the complementary of 5% circuit is more knowledgeable than the complementary of 50% circuit, indicating that the 5% complementary circuit likely possesses certain key parameters directly associated with knowledge retention, whereas the 50% complementary circuit may include a surplus of less critical parameters. In other words, a "theoretical circuit", if it exists, would incorporate more than 5% but less than 50% of parameters.
Table 2 provides additional evidence to how large the idealized circuits are. We replicate the experiments, this time extracting 15%, 25%, 35% parameters circuits while using the same hyperparameters. The table below shows the median post-edit negative log probabilities of the new target, combined with the results from the 50% and 5% circuits (higher values indicate worse performance). As expected, the edit performance of the extracted circuits declines as the circuit size decreases. If an idealized theoretical circuit exists that encapsulates all relevant knowledge in the model, it would likely have the worst edit performance due to strong ‘confirmation bias’. Therefore, we hypothesize that the 5% circuit may be closer to this idealized “theoretical circuit” than the 50% circuit. However, the exact size of this idealized circuit remains to be further investigated.
50% circuit 35% circuit 25% circuit 15% circuit 5% circuit
Circuit 7.78 8.10 8.28 9.30 9.37
Complementary circuit 0.0126 0.00999 0.00997 0.00985 0.00981
Table 2: Median post-Edit negative log probabilities of the new target across different percentage parameter circuits
Recent papers on circuit detection look for highly sparse circuits that usually contain less than 1% of the total parameters (Bayazit et al., 2023) or less than 3% of total edges (Hanna et al., 2024), depending on the tasks. Our observations indicate that a larger (i.e., >5%) subset of the networks on linguistic and knowledge-related tasks may also be worth of analysis.
Figure 3: Probability distribution of 5% circuit and 50% circuit. a. Post-edit probability of new target (top); b. Pre-edit probability of true target (bottom)
8 Do these circuits overlap?
The datasets used for circuit extraction in the previous sections showcase considerable diversity, yet they yield surprisingly consistent results. This consistency raises an interesting question: do these circuits significantly overlap?
Figure 4 features a heatmap that depicts the mask similarity across different tasks. Each entry in the heatmap represents the degree of overlap between 5% circuits from two datasets, measured relative to the base dataset listed in the row. For example, a value of ’0.137’ at the intersection of the second row and the first column suggests that the circuit derived from the ’dna1’ dataset overlaps with the ’h’ dataset circuit by 13.7% of the ’dna1’ circuit.
To create a baseline, we randomly select a 5% parameter circuit mask and calculate the overlap percentage of all parameters within the circuit. The results are presented in the table 3. The percentage is consistently around 5%.
dna1 dna2 h dnai1 dnawa1 dnawai1 e ra hhh
0.05026 0.05024 0.05013 0.05007 0.05021 0.05023 0.05001 0.04997 0.05007
Table 3: Percentage of parameter overlap between the circuits and a randomly selected circuit
The degrees of overlap between the circuits vary moderately across different dataset comparisons—indicating neither negligible nor overwhelming similarity. Interestingly, datasets that appear closely related, such as ’dna1’ and ’dna2’, sometimes display significant differences when viewed through the lens of the model’s mechanism. Notably, the ’ra’ (risk-averse) circuit shows greater overlaps with other circuits, suggesting that the risk-averse dataset may include terms that are broadly relevant across multiple contexts. Conversely, the ’e’ (extraversion) circuit demonstrates fewer overlaps, implying that the extraversion dataset likely features specialized terms that do not commonly occur in other datasets.
Additionally, we record the mask for the 50% circuit of 3 datasets and calculate the overlap percentage of all parameters within the circuit. Results are shown in Table 4. They each have approximately 60% overlap with one another, which is much higher than the overlap between the 5% circuits.
h circuit dna1 circuit ra circuit
h circuit 1 0.605 0.581
dna1 circuit 0.605 1 0.588
ra circuit 0.581 0.588 1
Table 4: Percentage of parameter overlap between 50% parameter circuits
Our finding echoes those of Merullo et al. (2023) that the circuits of different tasks may overlap with each other. Our analysis also underscores nuanced relationships between different datasets and highlights how specific dataset characteristics might influence the generalizability of the circuits derived from them.
Figure 4: Mask similarity across tasks heatmap. The overlaps are neither too large nor too small.
9 What Transformer modules are in the circuits?
In this section, we identify which types of layers are particularly influential within the extracted circuit. Table 5 presents a detailed breakdown of the proportion of total parameters that each type of layer accounts for within the circuit. The data reveal that LayerNorm constitutes a large percentage of the circuit’s parameters, especially noticeable within the 5% circuit subset. In contrast, the attentions and multi-layer perceptrons (MLPs) account for a smaller, yet comparable, share of the circuits.
This distribution of parameters is intriguing because it suggests that, while LayerNorm plays a crucial structural role, the computational responsibilities might be more evenly distributed or perhaps differently focused in attention and MLP layers. Notably, some studies found that the MLPs are critical repositories of knowledge within neural networks (Geva et al., 2020, 2022), and the attentions are critical for the mechanisms related to factual recall (Geva et al., 2023). Recently, Katz and Belinkov (2023) found that LayerNorm is related to the introduction of new tokens. There are also some recent debates about the exact localization of the knowledge (Hase et al., 2024; Niu et al., 2024). In contrast to prior works, our observations show that LayerNorm plays a critical role in the circuits.
Furthermore, the significant allocation of parameters to LayerNorm in the circuits raises questions about their specific function and impact on the overall network efficacy. Future analysis of LayerNorm could further advance our understanding.
dna1 5% circuit dna1 50% circuit h 5% circuit h 50% circuit
attn 0.04l 0.30 0.04 0.47
mlp 0.05 0.30 0.05 0.50
ln 0.37 0.63 0.37 0.62
Table 5: Proportion of total parameters accounted for by each layer type within the circuit
10 Discussion
Safety-related model behavior is relevant to the factual knowledge stored in the LLMs
The relationship between safety-related model behavior and the factual knowledge stored in large language models (LLMs) is intricate and significant. During the circuit extraction phase, we utilized model behavior datasets, such as those indicating ’risk aversion’ and ’extraversion’, in Model-Written Evaluations. These behavior-related circuits appear to possess a substantial amount of factual knowledge. Understanding LLM safety fundamentally involves comprehending model behavior. To achieve this, it is essential to understand the knowledge embedded within the model. However, the precise nature of the knowledge or concepts required for models to determine their behaviors remains an open question. Additionally, the mechanisms by which models use these concepts to guide their behavior are not fully understood. These areas represent critical open problems in the field.
Remove biases without compromising the model’s performance
One intriguing finding is that complementary circuits are easier to edit than the primary behavior-related circuits we identified. This suggests the possibility of performing editing tasks on complementary circuits without altering the primary circuits, thereby preserving the model’s behavior. If validated, this approach could offer a method to remove biases through editing tasks without compromising the model’s performance by using circuit identification.
Limitations and recommendations for future researchers
We propose several recommendations for future researchers aiming to enhance model safety: 1) Deeper Analysis of Complementary Circuits: Further investigation into the role and characteristics of complementary circuits can provide insights into how they interact with primary circuits and affect overall model behavior. 2) Searching for Theoretical Circuits: Investigating the possibility of theoretical circuits that can perform well for every specific tasks. 3) Utilizing Other Editing Techniques: Exploring other techniques like ROME (Meng et al., 2022a) to edit complementary circuits without impacting primary circuits may help mitigate biases and improve model fairness. 4) Expanding Behavior Datasets: Creating more comprehensive datasets that capture a wider range of behaviors can enhance our understanding of the relationships between behaviors and the knowledge stored in LLMs. By addressing these recommendations, future research can contribute to creating safer and more reliable language models.
11 Conclusion
We propose an approach to study the functions of the automatically-extracted circuits as a holistic object, through the lens of knowledge editing. By examining the effects of circuit size, the overlap between circuits derived from different datasets, and the distinct roles of various layers within these circuits, we present a comprehensive analysis that enhances our understanding of LLM behavior.
We observe a "confirmation bias" behavior at the circuit level: the primary circuits extracted from classification tasks, which are presumed to contain related knowledge, are challenging to edit. In contrast, complementary circuits, which are considered less integral to knowledge storage, adapt more readily to new information. This suggests that editing the primary circuits involves significant challenges, akin to altering well-established beliefs.
The impact of the circuit size was also explored. We compare the results of 5% circuits and 50% circuits, and reveal that larger circuits may introduce noise, diluting their effectiveness in knowledge interpretation and storage. This finding suggests that maybe essential knowledge is densely concentrated in an ideal "theoretical circuit", which is likely close to the 5% circuit.
Additionally, we investigate into the overlap between circuits from different linguistic task datasets, neural network behavior datasets and knowledge datasets. We reveals moderate variations in overlap degrees, suggesting a balance between distinctiveness and commonality among circuits.
Finally, the roles of different types of layers within circuits are examined, revealing that layer normalization layers command a large share of parameters, particularly in smaller circuits. Attention and multi-layer perceptron (MLP) layers hold smaller but critical shares. This distribution challenges traditional views on knowledge storage in neural networks and suggests that MLP layers, in particular, may be vital knowledge repositories. The prominence of layer normalization layers underscores their potential impact on network stability and performance, highlighting the need for deeper investigation into how these layers influence the overall efficacy of LLMs.
In conclusion, our research provides a detailed understanding of how automatically-extracted circuits within LLMs perform in knowledge editing tasks. These findings support future explorations into model behaviors, bias removal, and interpretability. By addressing the open problems and recommendations outlined in our discussion, future research can contribute to creating safer, fairer, and more reliable language models.
Appendix A 50% circuits probability distribution
This section presents the probability distributions for circuits discovered using different datasets, each representing 50% of the model’s parameters. The datasets used include BLiMP and Model-Written Evaluations. For this experiment, we utilized a GPU with 48GB of memory to perform the computations, with each dataset requiring less than one hour to process. The figures illustrate the post-edit negative log probability distributions of the new target distributions and the pre-edit probability distribution of the true target, providing insights into the impact of circuit-aware model editing.
Figures 5 and 6 display the post-edit probability of the new target and the pre-edit probability of the true target for the BLiMP dataset, respectively. Similarly, Figures 7 and 8 show these distributions for the Model-Written Evaluations dataset.
Overall, these visualizations provide a comprehensive view of the 50% circuits concerning knowledge editing.
Figure 5: Post-edit Negative Log Probability of New Target Distribution (50% circuit) (BLiMP).
Figure 6: Pre-edit Negative Log Probability of True Target Distribution (50% circuit) (BLiMP).
Figure 7: Post-edit Negative Log Probability of New Target Distribution (50% circuit) (Model-Written Evaluations).
Figure 8: Pre-edit Negative Log Probability of True Target Distribution (50% circuit) (Model-Written Evaluations).
Appendix B 5% circuits probability distribution
This section presents the probability distributions for circuits discovered using different datasets, which comprise 5% of the model’s parameters. The datasets used include BLiMP and Model-Written Evaluations. For this experiment, we also utilized a GPU with 64GB of memory to perform the computations, with each dataset requiring less than one hour to process. The figures illustrate the post-edit negative log probability distributions of the new target distributions and the pre-edit probability distribution of the true target, providing insights into the impact of circuit-aware model editing.
Figures 9 and 10 show the post-edit probability of the new target and the pre-edit probability of the true target for the BLiMP dataset, respectively. Figures 11 and 12 display the post-edit probability of the new target and the pre-edit probability of the true target for the Model-Written Evaluations dataset, respectively.
These visualizations offer a detailed perspective on the performance of the 5% circuits in the context of knowledge editing.
Figure 9: Post-edit Negative Log Probability of New Target Distribution (5% circuit) (BLiMP).
Figure 10: Pre-edit Negative Log Probability of True Target Distribution (5% circuit) (BLiMP).
Figure 11: Post-edit Negative Log Probability of New Target Distribution (5% circuit) (Model-Written Evaluations).
Figure 12: Pre-edit Negative Log Probability of True Target Distribution (5% circuit) (Model-Written Evaluations).